OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation x + 7y = -21. x intercept = (3, 0); y intercept = (0, 21) x intercept = (-3, 0); y intercept = (0, -21) x intercept = (21, 0); y intercept = (0, 3) x intercept = (-21, 0); y intercept = (0, -3)

OpenStudy (anonymous):

x intercept = (-21, 0); y intercept = (0, -3)

OpenStudy (anonymous):

is there anyway you could show me how you got that? i'm trying to figure out how to do it

OpenStudy (anonymous):

y intercept means when the graph crosses y axis, that happenes when x = 0 and opposite for x intercept where y = 0 knowing that, plug 0 in the equation, first time make x = 0 then solve for y the second time make y = 0 and solve for x x + 7y = -21 when x = 0, y = -21/7 = -3 when y= 0, x = -21